OpenStudy (anonymous):

Solve the following word problem by using an equation and factoring. The sum of a positive number and its square is 20. What is the number? Anwsers: 1, 2 , 3 , 4 , 5

OpenStudy (anonymous):

Well.. what seems to be the problem? :)

OpenStudy (anonymous):

i need the anwser..

OpenStudy (anonymous):

well what seems to be the answer? ;D

OpenStudy (anonymous):

I thought so... but.. why can't you figure it out yourself? Where are you stuck?

OpenStudy (anonymous):

I say is probably one of those in the delimited set of numbers that composed that group ;D

OpenStudy (anonymous):

x+x^2 = 20 x^2 + x - 20 = 0 Factor and solve the positive answer.

OpenStudy (anonymous):

\[\Large x+x^2=20 \] \[\Large x^2+x-20=0 \]

OpenStudy (anonymous):

wait don't tell me.... is a number who's has a pair of two primes, which is the lowest prime number there can be and is also a lowest pair number ;D

OpenStudy (anonymous):

is 4, ok?, just in case you be still wondering what number was it :p

OpenStudy (dumbcow):

its multiple choice so an easy trick is just try each possible answer , only 1 will work :) of course its best to be able to know how to solve quadratric equations without given answer choices

OpenStudy (anonymous):

NVM..